Neutron diffraction is used in determining the structures of molecules.

(a) Calculate the de Broglie wavelength of a neutron moving at 1.09% of the speed of light.
m

(b) Calculate the velocity of a neutron with a wavelength of 83 pm (1 pm = 10−12 m).
m/s
Answer #1

Solution-

using the formula p = mv

and we know that wavelength = h/mv

The mass of neutron = 1.67*10^-27
v = 0.0109*c

here c is the speed of light =3*10^8 m/s
= 6.626*10^-34/(1.67*10^-27*0.0109*3*10^8)
= 1.21*10^-13 m

b) v= h/m*λ

= 6.626*10^-34/(1.67*10^-27)(83*10^-12)
= 4.78*10^3 m/s